OpenStudy (rayman5766):

Variable y varies directly with variable x, and y = 6 when x = 20. What is the constant of variation in decimal form?

OpenStudy (rayman5766):

@cwrw238 help me flipprs

OpenStudy (anonymous):

Hi @rayman5766

OpenStudy (anonymous):

Whats your Possible Answers?

OpenStudy (anonymous):

Because.. If they vary directly then y = kx. Plug in the two values to find k. y = kx 6 = k(20) k = 6/20 = 3/10 = 0.3

OpenStudy (rayman5766):

oh ok

OpenStudy (anonymous):



OpenStudy (anonymous):

Got it? :)

OpenStudy (rayman5766):

so its 3 maby

OpenStudy (anonymous):